OpenStudy (anonymous):

2square root 2(radical 2-radical 10)+radical 3(radical 3+ radical15)

OpenStudy (anonymous):

\[2\sqrt{2}(\sqrt{2} - \sqrt{10}) +\sqrt{3}( \sqrt{3} + \sqrt{15})\]